40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 28 14:54:49 2016, ericq, Update, LSC, Status of the green PDH circuits 
    Reply  Thu Jan 28 19:10:16 2016, gautam, Update, LSC, Status of the green PDH circuits Servo_gain_calibration.pdf
Message ID: 11957     Entry time: Thu Jan 28 14:54:49 2016     Reply to this: 11958
Author: ericq 
Type: Update 
Category: LSC 
Subject: Status of the green PDH circuits 

Yesterday, I uploaded some EAGLE schematic files and a LISO source file for the green PDH servo electronics to the 40m LISO git repository. In doing so, I realized that the DCC document for the X box (D1400293) was not updated at the end of the electronics work we did in Aug/Sep 2014. This is entirely my fault. 

The Y box document (D1400294) is currently accurate. 

The missing information is that, as I posted In ELOG 10457, I ended up destroying our original X box, and replaced it with a spare from the ATF. It was restuffed to match the Y end box pretty much exactly. We will update the X circuit DCC page with an accurate schematic and photo. 

Gautam tells me that he and Rana were looking at the outdated schematic and thinking about improvements, but at least some of this was already done back in 2014 (specifically, the resistors used to specify the AD8336 preamp gain were changed).

ELOG V3.1.3-